Hardy and Littlewood conjectured that the the number of twin primes less than x is asymptotic to 2C_{2} ∫_{2}^{x} \frac{dt}{(log t)^{2}} where C_{2} is the twin prime constant. This has been shown to give excellent results for x up to about 10^{16}. This article presents statistics supporting the accuracy of the conjecture up to 10^{600}.